OpenStudy (anonymous):

Which of the following represents the absolute value of 7 - 5i?

hartnn (hartnn):

absolute value of a+bi is \(\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

isqr74

OpenStudy (calculator):

z=7 - 5i |z|= sqrt(7^2+5^2)=sqrt(74)

hartnn (hartnn):

yup, thats correct \(\sqrt{74}\)

OpenStudy (anonymous):

thank u it made it alot easier then the friken lesson

hartnn (hartnn):

welcome ^_^